OpenStudy (anonymous):

Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5? A. (1, 7) B. (5, 0) C. (2, -1) D. (0, 5)

OpenStudy (anonymous):

d. just plug in the answers to both eqs to see which answer yields a tru eq

OpenStudy (anonymous):

Yeah, but I wanna know the steps on how to do this.

OpenStudy (anonymous):

ok. y=-3x+5 y= 2x+5 so, -3x+5=2x+5 so, 5= 5x+5 so 0=5x so x=0 now plug in 0 for x in both eqs and see what you get

OpenStudy (anonymous):

Alright, I get it. Thanks! :D